AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sun Jun 8 11:55:00 2025
Author
*
:
Subject
*
:
<p><strong>12:52 </strong></p> <p><strong>Started Run 58 - Beam energy 200 keV/u.</strong></p> <p>Slow comparator threshold set to 0xc (130 keV) for all channels. Fast discriminators turned OFF for all channels by setting enable mask in discrinator control to 0xfffff. See attachments. </p> <p>D+ beam deuterium target. D+ beam energy is now 0.2 MeV/u. (beam calc calculation is 0.2037 MeV/u).</p> <p>All four detectors positioned fully out and stationary.</p> <p>Ring cycle time is 5 seconds.</p> <p>Cooler read value = 117.58 V, 0.00178 mA</p> <p>Target around <strong>8.9E12 </strong>atoms/cm2</p> <p>All checks ok!</p> <p>Beam on ionisation profile monitor is lost quite rapidly and looks to increase in size quite a lot due to target interactions. We may have to look at lowering target density</p> <p>Attachment 25 shows IPM with no target and attachment 26 shows IPM with the target 8e12 at/cm2. Compare also to attachment 12 (IPM with target at 2.6e12). </p> <p><strong>14:20</strong></p> <p>Checks are ok!</p> <p><strong>15:25</strong></p> <p>Fees 13-16 switched to MEC and bias on DSSD4 switched off</p> <p>All other checks ok</p> <p><strong>17:00</strong></p> <p>Target dropped to 1-2E12 atoms/cm2</p> <p>All checks okay- 17:25</p> <p><strong>17:45</strong></p> <p>Moved nozzle from x=-1500 to x=-2000, density went from 1E12 to 3E12</p> <p>Moved nozzle again at 18:00 from x=-2000 to x=-3000 in steps of 500, density at 4E12</p> <p><strong>18:45</strong></p> <p>Turned bias ON for dssd4 stable at V=10 but unstable at V=20 so bias turned back OFF</p> <p><strong>19:30</strong></p> <p>All checks ok!</p> <p>IPM attached, target density 2.6 atoms/cm2 (attachment 12)</p> <p><strong>20:00</strong></p> <p>Tried moving the nozzle more as target density dropped to 2-3E12 again</p> <p>Moved in 500 steps either side of x=-2500 but didn't improve, then moved in 100 steps from y=1600 to y=2300 and target density now 6E12 at x=-2500 and y=2300</p> <p>Tried dssd4 again but only stable to V=20 so bias OFF </p> <p><strong>21:50- </strong>All checks good, target density stable at 8E12</p> <p><strong>22:50-</strong> All checks fine (see attachments), target density at 6E12</p> <p><strong>00:10</strong></p> <p>Checks OK. Have zero'd statistics to see pause and resume statements over an hour.</p> <p>Rates looks OK. ~140 c/h 3He, and ~80 c/h for 3H and 1H each.</p> <p>Target has been stable at 7e12 at/cm2 for the last hour. </p> <p><strong>02:10</strong></p> <p>Checks OK. See pause statements after 2 hours - attachment 27, very few overall.</p> <p><strong>03:07</strong></p> <p>Checks OK. Target (9-11)E12</p> <p> </p> <p><strong>04:06</strong></p> <p>Checks OK. Target (9-11)E12</p> <p> </p> <p><strong>07:00</strong></p> <p>Checks OK. Target 9.8E12</p> <p>page <strong>AIDA/Check/Check.tml</strong>. failed</p> <p>ptoblem fixed</p> <p> </p> <p><strong>09:00</strong></p> <p>Checks OK. Target 8.9E12</p> <p> </p> <p><strong>10:32</strong></p> <p>No target data displayed for 10 minutes, but jet density unchanged. Restored display from remote access by M.L.</p> <p> </p> <p><strong>11:22</strong></p> <p>Storage data stopped and Faraday cup inserted to check the leakage current of Detector #4. Not good, current increases above 10 uA already at 40 V. </p> <p><strong>11:29</strong></p> <p>HV Detector #4 swithced off, Faraday cup removed, data stored again.</p> <p> </p> <p><strong>11:35</strong></p> <p>Check OK. Target 8.6E12</p> <p> </p> <p><strong>13:06</strong></p> <p>Check OK. Target 7.5E12</p> <p> </p> <p><strong>15:00</strong></p> <p>Temperature check - aida03 returns "No response". No response for rates either. Merger looks fine for 3? Likely just happened.</p> <p>Attempted to STOP the DAQ. All FEEs still going (refused to stop) and 3 in error. Decided to just turn all FEEs off.</p> <p>Restored DAQ with no issues. 3 is working. Temperatures fine. Rates fine. Unclear what had happened.</p> <p>Run59 is EMPTY. New run = run 60.</p> <p><strong>16:00, </strong>Full check, everything working after the restart</p> <p><strong>17:35</strong> Full check, everything is working smoothly!</p> <p><strong>18:35: </strong>Full check, status -Okay. thickness 5.7E12,</p> <p><strong>19:35:</strong> full ckeck! Status - OKAY, target theckness ~5E12</p> <p><strong>20:50 </strong>: Full check! status - okay</p> <p><strong>22:00 :</strong> Full check, status -okay. Target thickness 6.4E12</p> <p><strong>22:55</strong> : Full check, status -okay! ------ target thickness 5.8E12</p> <p><strong>01:00</strong> : Full check, all okay! </p> <p><strong>04:00</strong> : Full check, all okay! </p> <p><strong>07:00</strong> : Full check, all okay!</p> <p> </p> <p><strong>10:00</strong></p> <p>Attempted to change cooling to improve the beam. Changed cooling by 0.5V. Significant improvement in the IPM, but Rutherford goes down. Target missed?</p> <p>Changed orbit bumps. Increased Rutherford. IPM is back to how it was. Rates are possibly slightly higher? Will attempt more systematically in the afternoon, when we change beam energy again.</p> <p>Attempted to delay target on by 1 second. No significant changes in beam lifetime or rates. Can try systematically again in the afternoon.</p> <p>Restored settings to how they were.<strong> Started new run 61.</strong></p> <p>Temps OK</p> <p>Rates OK</p> <p>Merger OK</p> <p>Online histos look OK (did not reset)</p> <p> </p> <p>Checks</p> <p>Clock status</p> <p>FEE64 module aida16 global clocks failed, 6<br /> Clock status test result: Passed 15, Failed 1</p> <p>ADC calibration</p> <p>FEE64 module aida16 failed<br /> Calibration test result: Passed 15, Failed 1</p> <p>White Rabbit decoder</p> <p> <br /> Base Current Difference<br /> aida14 fault 0xbf08 : 0xbf08 : 0 <br /> aida14 : WR status 0x10<br /> White Rabbit error counter test result: Passed 15, Failed 1</p> <p>Understand the status reports as follows:-<br /> Status bit 3 : White Rabbit decoder detected an error in the received data<br /> Status bit 2 : Firmware registered WR error, no reload of Timestamp<br /> Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR</p> <p>FPGA Timestamp error counter test result: Passed 16, Failed 0</p> <p> </p> <p> </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Resubmit as new entry
Attachment 1:
Screenshot_from_2025-06-08_12-40-25.png
Original size: 1920x1080
Attachment 2:
Screenshot_from_2025-06-08_12-41-01.png
Original size: 1920x1080
Attachment 3:
Screenshot_from_2025-06-08_12-41-12.png
Original size: 1920x1080
Attachment 4:
Screenshot_from_2025-06-08_12-41-53.png
Original size: 1920x1080
Attachment 5:
Screenshot_from_2025-06-08_12-42-01.png
Original size: 1920x1080
Attachment 6:
Screenshot_from_2025-06-08_12-42-19.png
Original size: 754x506
Attachment 7:
Screenshot_from_2025-06-08_12-48-47.png
Original size: 1920x1080
Attachment 8:
Screenshot_from_2025-06-08_13-03-41.png
Original size: 1920x1080
Attachment 9:
Screenshot_from_2025-06-08_13-03-52.png
Original size: 1920x1080
Attachment 10:
Screenshot_from_2025-06-08_13-04-41.png
Original size: 1920x1080
Attachment 11:
Screenshot_from_2025-06-08_13-07-41.png
Original size: 1920x1080
Attachment 12:
IMG_4311.JPG
Original size: 1600x900
Attachment 13:
Screenshot_from_2025-06-08_14-12-02.png
Original size: 1711x909
Attachment 14:
Screenshot_from_2025-06-08_19-32-22.png
Original size: 1920x1080
Attachment 15:
Screenshot_from_2025-06-08_19-32-35.png
Original size: 1920x1080
Attachment 16:
Screenshot_from_2025-06-08_19-33-35.png
Original size: 1920x1080
Attachment 17:
Screenshot_from_2025-06-08_19-33-39.png
Original size: 1920x1080
Attachment 18:
Screenshot_from_2025-06-08_19-33-49.png
Original size: 1920x1080
Attachment 19:
Screenshot_from_2025-06-08_19-34-24.png
Original size: 1920x1080
Attachment 20:
Screenshot_from_2025-06-08_22-52-14.png
Original size: 1920x1080
Attachment 21:
Screenshot_from_2025-06-08_22-47-46.png
Original size: 1920x1080
Attachment 22:
Screenshot_from_2025-06-08_22-47-57.png
Original size: 1920x1080
Attachment 23:
Screenshot_from_2025-06-08_22-48-27.png
Original size: 1920x1080
Attachment 24:
Screenshot_from_2025-06-08_22-48-36.png
Original size: 1920x1080
Attachment 25:
20250608_121153.jpg
Original size: 4000x3000
Attachment 26:
20250608_124137.jpg
Original size: 4000x3000
Attachment 27:
Screenshot_from_2025-06-09_02-06-22.png
Original size: 1920x1014
Attachment 28:
Screenshot_from_2025-06-09_02-06-22.png
Original size: 1920x1014
Attachment 29:
Screenshot_from_2025-06-09_03-05-16.png
Original size: 1920x1080
Attachment 30:
Screenshot_from_2025-06-09_03-06-47.png
Original size: 1920x1080
Attachment 31:
Screenshot_from_2025-06-09_15-51-46.png
Original size: 1920x1080
Attachment 32:
Screenshot_from_2025-06-09_15-52-16.png
Original size: 1920x1080
Attachment 33:
Screenshot_from_2025-06-09_15-52-50.png
Original size: 1920x1080
Attachment 34:
Screenshot_from_2025-06-09_15-53-06.png
Original size: 1920x1080
Attachment 35:
Screenshot_from_2025-06-09_15-53-22.png
Original size: 1920x1080
Attachment 36:
Screenshot_from_2025-06-09_15-55-24.png
Original size: 1920x1080
Attachment 37:
Screenshot_from_2025-06-09_15-56-00.png
Original size: 1920x1080
Attachment 38:
Screenshot_from_2025-06-09_15-56-19.png
Original size: 1920x1080
Attachment 39:
Screenshot_from_2025-06-09_15-57-26.png
Original size: 1920x1080
Attachment 40:
Screenshot_from_2025-06-09_15-58-04.png
Original size: 1920x1080
Attachment 41:
Screenshot_from_2025-06-09_21-54-32.png
Original size: 1920x1080
Attachment 42:
Screenshot_from_2025-06-09_21-55-03.png
Original size: 1920x1080
Attachment 43:
Screenshot_from_2025-06-09_21-55-21.png
Original size: 1920x1080
Attachment 44:
Screenshot_from_2025-06-09_21-55-30.png
Original size: 1920x1080
Attachment 45:
Screenshot_from_2025-06-09_21-55-41.png
Original size: 1920x1080
Attachment 46:
Screenshot_from_2025-06-09_21-55-49.png
Original size: 1920x1080
Attachment 47:
Screenshot_from_2025-06-10_10-05-08.png
Original size: 1920x1080
Attachment 48:
Screenshot_from_2025-06-10_10-05-29.png
Original size: 1920x1080
Attachment 49:
Screenshot_from_2025-06-10_10-05-39.png
Original size: 1920x1080
Attachment 50:
beamCalc-0.2.png
Original size: 276x835
Maximum number of attachments reached
Draft saved at 00:00:00
ELOG V3.1.3-7933898